A tank is 5/8 full, 180 litres have left the tank and is now 1/6 full. How many litres are left in the tank?
Answer by rfer(16322)

15/24-4/24=11/24 gone
180/(11/24)=180*24/11=393 litres total
393*1/6=393/6=65.5 litres left